OpenStudy (anonymous):

Factor x^2 - 8x + 9 :/

OpenStudy (anonymous):

:/ what's it mean

OpenStudy (campbell_st):

I don't think if can be factored

OpenStudy (campbell_st):

my solution to your problem is general quadratic formula \[x = \frac{8 \pm \sqrt{(-8)^2 - 4 \times1\times9}}{2\times1}... = \frac{8 \pm \sqrt{28}}{2}\]\\ or \[x = 4 \pm \sqrt{7}\]

OpenStudy (anonymous):

that's ur answer is right???

OpenStudy (campbell_st):

no if its factorised its \[(x + 4 + \sqrt{7})(x + 4-\sqrt{7})\]